40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 25 18:20:15 2015, ericq, Update, LSC, ALS reconstruction in progress 
    Reply  Fri Jun 26 03:24:18 2015, ericq, Update, LSC, ALS reconstruction in progress 
       Reply  Mon Jun 29 12:28:45 2015, ericq, Update, LSC, ALS reconstruction in progress 
          Reply  Tue Jun 30 05:47:38 2015, rana, Update, LSC, Use BALUNs 
Message ID: 11379     Entry time: Fri Jun 26 03:24:18 2015     In reply to: 11378     Reply to this: 11381
Author: ericq 
Type: Update 
Category: LSC 
Subject: ALS reconstruction in progress 

Too sleepy to make full ELOG. Stay tuned. 

Two 25dB amplifiers (with fins!) are living in the top shelf on the PSL table, inputs currently grounded. I broke out the fused 24V power from the AOM driver to power the two amps and the AOM driver. I used the POP55 and AS165 heliax cables to get their outputs to the LSC rack, through delay lines, into demod board. 

Driving with -20dBm at 55MHz, the BEATX signal chain has about 60Hz RMS noise, which is about what I measured for driving the old beatbox with a marconi. High frequency noise is a much nicer shape, though. The BEATY signal didn't seem to be getting through, will double check soon. 

Still old delay cables, not nicely shielded or isolated or anything. We'll have to pipe the monitor signal from the LSC rack over to the control room analyzer now. 

 

ELOG V3.1.3-